Two planes are being evaluated, with the projected life of each being 10 years. The cash flow for each plane are below. Use MARR of 15%. Based on the rate of return which is the most better.

A B
First Cost $350 000 $415 000
M & O Costs 15 000 10 000
Annual Benefit 90 000 125 000
Salvage Value 35 000 45 000
Answer #1
A B
First Cost 350000 415000
M & O Costs 15 000 10 000
Annual Benefit 90 000 125 000
Salvage Value 35 000 45 000
IRR is that discount rate which will give 0 NPV, that is for which
PV of cash inflows = PV of cash outflows.
The rate has to be found out by trial and error, such that
NPV = 0
IRR for Plane A:
For NPV = 0, the equation would be
0 = -350000+(90000-15000)*PVIFA(r,10)+35000*PVIF(r,10)
where r = IRR
The value of r is to be found out by trial and error using different discount rates.
Using 18% as discount rate
NPV = -350000+75000*4.49409+35000*0.19106 = -6256.15
Using 17% as discount rate
NPV = -350000+75000*4.65860+35000*0.20804 = 6676.40
IRR lies between 17% and 18%. The value of r can be found out by simple
interpolation as below:
=17+6676.40/(6256.15+6676.40) = 17.52%
IRR for Plane B:
Using discount rate of 25%, NPV =
= -415000+115000*3.57050+45000*0.10737 = 439.15
Using discount rate of 26%, NPV =
= -415000+115000*3.46481+45000*0.09915 = -12085.10
IRR = 25%+439.15/(12085+439.15)/100 = 25.04%
DECISION:
Both the alternatives are acceptable as their IRR is greater than the MARR.
But, as the IRR of Plane B is more it is the better choice.